### v1

@@ 3.3 Parametric Method: Propensity Score Matching (PSM)
$\text{logit}\Big{(}P(\text{Gender}_{i}=1\mid X_{i})\Big{)}=$ $\beta_{0}+\sum_{j}\beta_{1j}\,(\text{University Code}_{ij}\times\text{Department Code}_{ij}\times\log_{10}(\text{i10index}_{i}))$
$+\sum_{k}\beta_{2k}\,(\text{Titles}_{ik}\times\text{Working Years}_{i})$
$+\beta_{3}\,(\text{Google Scholar ID}_{i}),$
where $\text{logit}(p)=\ln\left(\frac{p}{1-p}\right)$ and $p=P(\text{Gender}=1\mid X)$. Here, each $\beta_{1j}$ and $\beta_{2k}$ corresponds to a dummy-coded interaction term, reflecting the fact that both the university affiliation and academic title variables are represented using indicator variables in the model. The first captures the joint effects of institutional affiliation and research productivity, while the second reflects the interaction between academic rank and experience. The binary Google Scholar ID variable indicates academic presence and visibility. This variable plays a greater role in the full dataset, where it distinguishes between entries with and without a visible publication record; in the reduced dataset, it is constant and contributes little variation. Interaction terms were selected through an iterative model development process guided by balance diagnostics, including standardized mean differences and covariate balance plots. Simpler models without these interactions resulted in poorer balance and weaker covariate overlap, underscoring the importance of capturing these complex structures in academic settings.
After estimating the propensity scores, we performed nearest-neighbor matching without replacement, using a caliper of 0.2 standard deviations of the logit of the propensity score. Post-matching diagnostics confirmed improved covariate balance across treatment groups.
To estimate the causal effect of gender on salary, we subsequently applied an ordinary least squares (OLS) regression to the matched sample. The regression model included the same set of covariates and interaction terms as those used in the propensity score model, ensuring consistent adjustment for confounding structures. Specifically, the model is formulated as:
$\log(\text{Salary}_{i})=$ $\alpha+\tau\cdot\text{Gender}_{i}$
@@ 3.4 Non-Parametric Method: Causal Forest
To model complex relationships and allow for treatment effect heterogeneity, we employed a non-parametric method known as the Causal Forest, developed by Athey and Imbens (2016) and extended by Wager and Athey (2018). Unlike traditional regression-based approaches, which impose parametric assumptions and often yield only population-level average treatment effects, the causal forest estimates the Conditional Average Treatment Effect (CATE), denoted as $\tau(X_{i})$, for each unit based on its observed covariates:
@@
where $\hat{\tau}_{t}(X_{i})$ is the treatment effect estimate from the $t$-th tree in the ensemble.
Causal forests are built upon the generalized random forest framework and introduce several key innovations adjusted for causal inference. First, the algorithm adopts an honest estimation strategy through sample-splitting: one portion of the data is used to determine tree structure (i.e., splitting rules), and a separate, disjoint portion is used to estimate treatment effects within each leaf. This reduces overfitting and improves the validity of inference. Second, the method incorporates orthogonalization, also known as double machine learning, to reduce sensitivity to nuisance parameter estimation. In practice, this means that the model first estimates the propensity score and the conditional expectation of the outcome, then fits the forest using residualized versions of the treatment and outcome variables. Third, treatment effect estimation is localized—each tree partition defines neighborhoods of similar observations, and the treatment effect is estimated by comparing outcomes between treated and control units within these local regions. This enables the method to flexibly model high-order, nonlinear interactions without the need for manual specification.
In our analysis, the covariate matrix $X_{i}$ included academic title, department code, university code, years of experience, log-transformed i10-index, and Google Scholar ID. These covariates capture essential dimensions of professional background, institutional affiliation, and research productivity, and align with those used in the parametric model for consistency across approaches.
@@ 4.1 Non-Causal Analysis
The Ordinary Least Squares (OLS) regression results, presented in Table C.2, indicate that even after adjusting for these covariates, a statistically significant gender pay gap persists. The estimated coefficient for gender is $-0.0321$ ($p=2.32e^{-16}$), suggesting that female faculty members earn approximately 7.12% less than male faculty members with similar observed characteristics.
@@ 4.2 Parametric Causal Analysis: Propensity Score Matching (PSM) Model
We evaluated the matching process using numerical and visual methods, ensuring well-balanced treatment and control groups. The sample consists of 3,062 male and 1,920 female faculty members, with 1,919 successful matches. The balance diagnostics confirm that post-matching standardized mean differences (SMD) are near zero for all key covariates, ensuring comparability between treatment and control groups. The full covariate balance table is available in Table 4. As previously demonstrated, the Love Plot (Figure 5) confirms that propensity score matching successfully balances key covariates across gender groups.
We then performed a linear regression on the matched data (see Table C.3 for full results). Gender remains a significant predictor of salary, with a coefficient of $\beta=-0.0297$ (SE $=0.0054$, $p=3.61e^{-8}$), indicating that female faculty earn approximately 6.61% less than their male counterparts with similar backgrounds. This confirms a persistent gender wage gap, even after accounting for institutional, departmental, and career-related factors, emphasizing the need for targeted interventions in academic salary structures. Our results also showed that career stage, research productivity, and disciplinary context jointly shape salary.
@@ 4.3 Non-Parametric Causal Estimation: Causal Forest
Using the Causal Forest model as a non-parametric causal inference method, we analyzed both overall trends and individual variations in gender-based salary disparities.
With Google Scholar ID subset, the estimated Average Treatment Effect (ATE) was $-0.0260$ (SE = $0.0039$), indicating that, on average, female faculty earn 5.81% less than male faculty after adjusting for covariates. This confirms a persistent gender wage gap even after controlling for critical confounding variables.
To explore individual variations, we estimated Individual Treatment Effects (ITEs) using the causal forest model. The distribution of ITEs, as shown in Appendix C.4.1, reveals substantial heterogeneity in the gender effect on salaries. Estimated treatment effects range from $-0.14$ to $0.04$.
[…]
The findings from the causal forest model confirm the existence of a gender-based salary disparity in academia, consistent with the results concluded in PSM analysis. However, observed differences underscore the importance of considering individual characteristics such as tenure and research productivity.
@@
Full Dataset -0.0293 6.52%
Causal Forest Google Scholar Subset -0.0252 5.64%
Full Dataset -0.0266 5.94%
@@ 5 Conclusion and Discussion
In this study, we investigated the factors that might lead to the gender gap in faculty salary by studying approximately 12 thousand faculty affiliated to 16 universities in the University of North Carolina system from year 2022. Our results reveal persistent salary disparities across all career stages, academic performances, institutional types, and department types.
The raw wage gap shows female faculty earning 11.71-12.80% less than their male counterparts. Even after controlling for various factors through propensity score matching and causal forest analysis, significant disparities of 5.64-6.61% remain, indicating structural inequities in academic compensation. The impact of gender on salary varies obviously across career stages. Early-career faculty face a consistent negative effect of approximately 5.59%, while mid-career stages show increased volatility in wage differences. Senior faculty members still experience a smaller but persistent gap around 5.53%, suggesting that gender-based disparities last throughout academic careers. Furthermore, growth in academic performances does not eliminate the gender wage gap, shown by the persistent negative effects across all research impact levels. The U-shaped relationship, with the largest gender gaps occurring at moderate publication impact levels, reflects that the worst case are found among faculty with mid-range academic performance. Then, institutional context plays a crucial role, with doctoral universities showing stronger positive relationships between research productivity and salary, while bachelor’s and master’s institutions demonstrate weaker or negative relationships. Lastly, departmental variations emerge as another significant factor. Business schools show the strongest positive returns for research productivity, while arts and humanities consistently show negative relationships, perhaps indicating that Google Scholar does not properly measure scholarly productivity in these fields. STEM fields exhibit varying patterns depending on institution type, highlighting the complex interaction between department, institution, and gender in determining faculty compensation. These findings emphasize the need for comprehensive policy reforms in academic compensation. Institutions should implement transparent pay structures, standardize research output evaluation processes, and conduct regular equity audits across departments. Additionally, particular attention should be paid to addressing early-career wage gaps and reviewing promotion criteria and timing.
There are several limitation of this work. First, the reliance on Google Scholar metrics as a measure of academic productivity excludes other important aspects of faculty contribution such as teaching excellence, service commitments, and mentorship activities (Wildgaard et al., 2014). Then, the cross-sectional nature of our 2022 data limits our ability to examine how salary disparities change and evolve throughout individuals’ careers. Additionally, the focus on the UNC system, while providing a complete institutional framework, may limit the generalizability of findings to private institutions or other state systems with different governance structures and policies (Kuhn and Johnson, 2019). Future research would benefit from longitudinal studies investigating how these disparities evolve over time, as well as investigations into the intersection of gender with demographic factors.
@@ A.1 Distribution of Gender Assignment Methods
Automated Assignment ($\geq$60confidence) 11505 95.56
@@
      Names with confidence $<$60% 274 2.28
@@ B.3 Consistency: Treatment Definition and No Interference
Titles (Associate Professor) 0.0526 0.0049 $<$2e-16 ***
Titles (Professor) 0.1760 0.0057 $<$2e-16 ***
@@ C.2 Non-Causal Analysis
Titles: Associate Professor 0.0526 0.0049 10.750 $<$2e-16***
Titles: Professor 0.1760 0.0057 30.753 $<$2e-16***
University Code (DRU(H)) 0.0567 0.0067 8.511 $<$2e-16***
University Code (DU/VA) 0.1693 0.0066 25.501 $<$2e-16***
Department Code (B) 0.2743 0.0083 33.198 $<$2e-16***
Department Code (MHS) 0.2674 0.0067 39.647 $<$2e-16***
Department Code (NS) 0.0754 0.0066 11.404 $<$2e-16***
Department Code (SS) 0.0748 0.0070 10.727 $<$2e-16***
Department Code (TE) 0.0971 0.0066 14.714 $<$2e-16***
@@ C.3 Parametric Causal Analysis
Titles: Assistant Professor $\times$ Working Years -0.0083 0.0007 -12.075 $<$2e-16***
@@
Titles: Professor $\times$ Working Years 0.0030 0.0003 9.215 $<$2e-16***
@@
DRU(H) $\times$ Dept. AH $\times$ log10(i10-index) -0.0828 0.0086 -9.671 $<$2e-16***
@@
DRU(H) $\times$ Dept. B $\times$ log10(i10-index) 0.0938 0.0113 8.314 $<$2e-16***
DU/VA $\times$ Dept. B $\times$ log10(i10-index) 0.1610 0.0121 13.323 $<$2e-16***
@@
DU/VA $\times$ Dept. MHS $\times$ log10(i10-index) 0.1479 0.0056 26.450 $<$2e-16***

### v2

@@ 3.3 Parametric Method: Propensity Score Matching (PSM)
$\text{logit}\Big(P(\text{Gender}_{i}=1\mid X_{i})\Big)=\$$\beta_{0}+\sum_{j}\beta_{1j}\,(\text{University Code}_{ij}\times\text{Department Code}_{ij}\times\log_{10}(\text{i10index}_{i}))$
$+\sum_{k}\beta_{2k}\,(\text{Titles}_{ik}\times\text{Working Years}_{i}),$
where $\text{logit}(p)=\ln\left(\frac{p}{1-p}\right)$ and $p=P(\text{Gender}=1\mid X)$. Here, each $\beta_{1j}$ and $\beta_{2k}$ corresponds to a dummy-coded interaction term, reflecting the fact that both the university affiliation and academic title variables are represented using indicator variables in the model. The first captures the joint effects of institutional affiliation and research productivity, while the second reflects the interaction between academic rank and experience. Although Google Scholar ID is an indicator of academic presence and visibility, including it in the propensity score model substantially worsened post-matching covariate balance. Therefore, in the full dataset we excluded Google Scholar ID from the propensity score process to improve balance, and instead incorporated it in the outcome regression as an adjustment covariate. In the reduced subset, Google Scholar ID is constant and thus plays no role in the analysis. So, wo did not include Google Scholar ID in the propensity score process. Interaction terms were selected through an iterative model development process guided by balance diagnostics, including standardized mean differences and covariate balance plots. Simpler models without these interactions resulted in poorer balance and weaker covariate overlap, underscoring the importance of capturing these complex structures in academic settings.
After estimating the propensity scores, we performed nearest-neighbor matching without replacement, using a caliper of 0.2 standard deviations of the logit of the propensity score. We chose nearest-neighbor matching because it is a widely used and interpretable approach that pairs treated and control units with the closest estimated propensity scores, thereby producing samples that are easy to compare directly. The caliper size of 0.2 was selected following common practice in the causal inference literature, where simulation studies have shown that this threshold effectively reduces bias without discarding excessive observations (austin2011optimal). Post-matching diagnostics confirmed improved covariate balance across treatment groups.
To estimate the causal effect of gender on salary, we subsequently applied an ordinary least squares (OLS) regression to the matched sample. Unlike the propensity score process, here we included Google Scholar ID as a covariate in full dataset, along with the same set of interactions, to adjust for publication visibility. Specifically, the model is formulated as:
$\log(\text{Salary}_{i})=\$$\alpha+\tau\cdot\text{Gender}_{i}$
@@ 3.4 Non-Parametric Method: Causal Forest
To model complex relationships and allow for treatment effect heterogeneity, we employed a non-parametric method known as the Causal Forest, developed by athey2016recursive and extended by wager2018estimation. Unlike traditional regression-based approaches, which impose parametric assumptions and typically focus on population-level average treatment effects, causal forests estimate the Conditional Average Treatment Effect (CATE), denoted as $\tau(X_{i})$, for each unit based on its observed covariates. A natural way to express this is as an ensemble average across trees:
@@
where $\hat{\tau}_{t}(X_{i})$ is the treatment effect estimate from the $t$-th tree. This formulation emphasizes that the CATE is a point estimate conditional on covariates $X_{i}$, obtained as a weighted average across trees.
In practice, however, causal forests are not simply tree averages. Following the generalized random forest framework (athey2019estimating), they are implemented as an orthogonalized local CATE estimator:
$\hat{\tau}(x)=\frac{\sum_{i=1}^{n}\alpha_{i}(x)\,\big(Y_{i}-\hat{m}^{(-i)}(X_{i})\big)\,\big(Z_{i}-\hat{e}^{(-i)}(X_{i})\big)}{\sum_{i=1}^{n}\alpha_{i}(x)\,\big(Z_{i}-\hat{e}^{(-i)}(X_{i})\big)^{2}},$
where $\alpha_{i}(x)$ are adaptive weights determined by the forest structure and can be interpreted as a data-driven kernel, $\hat{m}^{(-i)}(X_{i})$ is the estimated conditional mean outcome and learned from a separate subsample, and $\hat{e}^{(-i)}(X_{i})$ is the estimated propensity score. The superscript (i)(-i) indicates that the estimates are obtained without observation $i$, which prevents overfitting. This orthogonalization, also known as double machine learning, step reduces overfitting and bias from nuisance parameter estimation.
Beyond this estimator, causal forests incorporate several algorithmic innovations adjusted for causal inference. First, the algorithm adopts an honest estimation strategy through sample-splitting: one portion of the data is used to determine tree structure (i.e., splitting rules), and a separate, disjoint portion is used to estimate treatment effects within each leaf. This reduces overfitting and improves the validity of inference. Second, the method incorporates orthogonalization, also known as double machine learning, to reduce sensitivity to nuisance parameter estimation. In practice, this means that the model first estimates the propensity score and the conditional expectation of the outcome, then fits the forest using residualized versions of the treatment and outcome variables. Third, treatment effect estimation is localized—each tree partition defines neighborhoods of similar observations, and the treatment effect is estimated by comparing outcomes between treated and control units within these local regions. This enables the method to flexibly model high-order, nonlinear interactions without the need for manual specification.
In our analysis, the covariate matrix $X_{i}$ included academic title, department code, university code, years of experience, and log-transformed i10-index. For the subset of faculty with valid Google Scholar IDs, we used this specification directly, without including the ID variable. For the full sample, however, we aligned the nuisance models with the parametric framework: the propensity nuisance $\hat{e}(X)$ was estimated without Google Scholar ID, while the outcome nuisance $\hat{m}(X,Z)$ incorporated it as a binary indicator. This ensured comparability across approaches and improved covariate balance, since including Google Scholar ID in the propensity score stage led a less balanced post-matching.
We implemented causal forests using the grf package in R (athey2019generalized). For the subset sample, we grew 3,000 trees to stabilize estimates, used honest splitting to reduce bias, and set the minimum node size to 5 and the subsample fraction to 0.5, following common practice in causal forest estimation. We tuned only the number of covariates considered at each split (mtrymtry), which balances predictive accuracy and computational efficiency. For the full sample, all major hyperparameters were tuned automatically to accommodate the higher-dimensional nuisance functions. To address extreme estimated propensities close to 0 or 1, we focused on the average treatment effect for the overlap population, which emphasizes comparisons among faculty with comparable treatment probabilities and yields more reliable inference than the full-sample ATE.
@@ 4.1 Non-Causal Analysis
The Ordinary Least Squares (OLS) regression results, presented in Table C.2, indicate that even after adjusting for these covariates, a statistically significant gender pay gap persists. The estimated coefficient for gender is 0.0321-0.0321 ($p=2.32e^{-16}$), suggesting that female faculty members earn approximately 7.12% less than male faculty members with similar observed characteristics.
@@ 4.2 Parametric Causal Analysis: Propensity Score Matching (PSM) Model
We evaluated the matching process using numerical and visual methods, ensuring well-balanced treatment and control groups. The sample consists of 3,062 male and 1,920 female faculty members. After matching, 1,919 females were successfully paired with 1,179 comparable males, leaving 1,883 males and 1 female unmatched. The balance diagnostics confirm that post-matching standardized mean differences (SMD) are near zero for all key covariates, ensuring comparability between treatment and control groups. The full covariate balance table is available in Table 4. As previously demonstrated, the Love Plot (Figure 5) confirms that propensity score matching successfully balances key covariates across gender groups.
We then performed a linear regression on the matched data (see Table C.3 for full results). Gender remains a significant predictor of salary, with a coefficient of $\beta=-0.0297$ (SE =0.0054=0.0054, $p=3.61e^{-8}$), indicating that female faculty earn approximately 6.61% less than their male counterparts with similar backgrounds. This confirms a persistent gender wage gap, even after accounting for institutional, departmental, and career-related factors, emphasizing the need for targeted interventions in academic salary structures. Our results also showed that career stage, research productivity, and disciplinary context jointly shape salary.
@@ 4.3 Non-Parametric Causal Estimation: Causal Forest
Using the Causal Forest model as a non-parametric causal inference method, we analyzed both overall trends and individual variations in gender-based salary disparities. The validity of these estimates depends on the assumption that, conditional on observed covariates, including rank, department, institutional affiliation, working years, and research productivity, gender assignment is unconfounded. The honesty principle embedded in causal forests helps reduce overfitting and bias in treatment effect estimation by separating model training from effect prediction (wager2018estimation).
For the subset of faculty with valid Google Scholar IDs, the estimated Average Treatment Effect (ATE) for the overlap population was 0.0267-0.0267 (SE = $0.0039$), indicating that, among comparable faculty, female faculty earn on average 5.96% less than male faculty after adjusting for covariates. This confirms a persistent gender wage gap even after controlling for critical confounding variables. Sensitivity checks varying key hyperparameters yielded substantively unchanged results (see Appendix C.4.1).
To explore individual-level variation, we estimated Individual Treatment Effects (ITEs) using the causal forest model. The distribution of ITEs, shown in Appendix C.4.2, ranges from 0.035-0.035 to 0.021-0.021, underscoring that the gender effect on salaries is far from uniform. Such heterogeneity indicates that average estimates may conceal important subgroup differences, motivating a closer examination of where disparities are most obvious.
We visualized ITEs by coloring observations with different covariates to detect systematic patterns. Among the explored factors, department and research productivity revealed the clearest structure, with Medical and Health Sciences (MHS) standing out as the field where gender inequities are particularly acute.
Figure C.4.3 displays ITEs by career stage across departments. In most fields, the gender gap remains relatively stable over the career path, but MHS faculty show a consistently larger salary disadvantage for women. On average, women in MHS earn about 7% less than comparable men, compared to least gaps around 5.5% in Natural and Social Sciences (NS, SS) and about 6% in Arts and Humanities (AH), Business (B), and Technology and Engineering (TE). The separation between MHS and other disciplines widens with experience, suggesting persistent structural barriers in this field.
Figure C.4.4 conditions on research productivity. For MHS faculty, the estimated penalty starts around 6% at lower productivity levels, worsens to more than 7% in the mid-range, and then shows slight improvement at the highest levels, though still remaining near 7%. By contrast, other departments display relatively little variation, with gaps largely concentrated between 5–6%. This pattern suggests that women in MHS face a particularly persistent and pronounced disadvantage, especially at intermediate levels of research productivity, whereas disparities in other fields are smaller and more stable.
These findings reinforce the view that structural disparities in faculty compensation are not merely average effects but reflect deeper patterns shaped by tenure, institutional norms, and how scholarly output is translated into financial reward. Non-parametric causal estimation thus provides a flexible tool for identifying when and where inequities are most acute, offering valuable insights for both empirical researchers and academic administrators.
@@
Full Dataset -0.0269 6.01%
Causal Forest Google Scholar Subset -0.0267 5.96%
Full Dataset -0.0262 5.99%
@@ 5 Conclusion and Discussion
Descriptive comparisons reveal that, on average, female faculty earn 11.71% to 12.80% less than male colleagues. Even after adjusting for institutional, disciplinary, and productivity-related factors using causal inference methods, substantial gaps remain: propensity score matching yields a disparity of 6.61%, and causal forests estimate an average treatment effect of 5.96%. These results provide obvious evidence of structural gender-based salary inequities across the North Carolina public university system.
The gender wage gap is not uniform. Heterogeneity analyses show that disparities vary systematically across departments, with Medical and Health Sciences (MHS) standing out as the field where inequities are most acute. Female faculty in MHS earn approximately 7% less than comparable men, a penalty that persists across career stages and remains visible even when conditioning on research productivity. By contrast, gaps in Natural Sciences and Social Sciences are closer to 5.5%, while Arts and Humanities, Business, and Technology and Engineering average around 6%. This suggests that women in MHS face more persistent structural barriers to pay equity, whereas disparities in other fields, while present, are less severe.
Institutional and disciplinary contexts also shape compensation patterns. Business and Medical and Health Sciences departments, particularly at R1 (DU/VA) institutions, demonstrate the strongest positive returns to research productivity. In contrast, the Arts and Humanities exhibit consistently negative associations between salary and citation-based metrics, raising concerns about the suitability of bibliometric indicators in evaluating scholarly impact in these fields. Other fields (Technology and Engineering, Natural Sciences, Social Sciences) display mixed patterns: research output is rewarded at R1 institutions but often penalized or ignored at less research-intensive settings. These findings underscore the interaction between institution type, academic discipline, and evaluative metrics in determining pay structures.
Taken together, our findings highlight the need for comprehensive policy reforms. Institutions should improve salary transparency, standardize criteria for evaluating research output across disciplines, and conduct regular equity audits disaggregated by career stage, institution type, and academic field. Targeted interventions, such as early-career mentoring, off-scale salary review processes, and promotion pathway monitoring, may help address disparities before they become entrenched.
This study has several limitations. First, academic productivity is measured solely through Google Scholar, which may underrepresent contributions such as teaching excellence, mentorship, and institutional service (ACE2022equity). Second, the analysis is cross-sectional and does not capture how disparities evolve over time. Third, while the UNC system provides a coherent institutional framework, the findings may not generalize to private universities or public systems with different governance structures (kuhn2019feature). Future work would benefit from longitudinal analyses, broader definitions of scholarly impact, and investigation of intersectional factors, such as race, marital status, and contract type, that may further shape pay equity in academia.
@@ A.1 Distribution of Gender Assignment Methods
Automated Assignment ($\geq$60% confidence) 11505 95.56
@@
    Names with confidence <<60% 274 2.28
@@ B.3 Consistency: Treatment Definition and No Interference
Titles (Associate Professor) 0.0526 0.0049 <<2e-16 ***
Titles (Professor) 0.1760 0.0057 <<2e-16 ***
@@ C.2 Non-Causal Analysis
Titles: Associate Professor 0.0526 0.0049 10.750 <<2e-16***
Titles: Professor 0.1760 0.0057 30.753 <<2e-16***
University Code (DRU(H)) 0.0567 0.0067 8.511 <<2e-16***
University Code (DU/VA) 0.1693 0.0066 25.501 <<2e-16***
Department Code (B) 0.2743 0.0083 33.198 <<2e-16***
Department Code (MHS) 0.2674 0.0067 39.647 <<2e-16***
Department Code (NS) 0.0754 0.0066 11.404 <<2e-16***
Department Code (SS) 0.0748 0.0070 10.727 <<2e-16***
Department Code (TE) 0.0971 0.0066 14.714 <<2e-16***
@@ C.3 Parametric Causal Analysis
Titles: Assistant Professor $\times$ Working Years -0.0083 0.0007 -12.075 <<2e-16***
@@
Titles: Professor $\times$ Working Years 0.0030 0.0003 9.215 <<2e-16***
@@
DRU(H) $\times$ Dept. AH $\times$ log10(i10-index) -0.0828 0.0086 -9.671 <<2e-16***
@@
DRU(H) $\times$ Dept. B $\times$ log10(i10-index) 0.0938 0.0113 8.314 <<2e-16***
DU/VA $\times$ Dept. B $\times$ log10(i10-index) 0.1610 0.0121 13.323 <<2e-16***
@@
DU/VA $\times$ Dept. MHS $\times$ log10(i10-index) 0.1479 0.0056 26.450 <<2e-16***